Why corporation issued 200,000 shares of $20 par value, 6% preferred stock on January 1, 2017 for $4,500,000. In December 2018, Eby declared its first dividend of $800,000. If the preferred stock is not cumulative, how much of the $800,000 would be paid to common stockholders? Of the preferred stock is cumulative, how much of the $800,000 would be paid to common stockholders.
Answer #1

Calculate following

Annual preferred stock = 200000*20*6% = 240000

When preferred stock is not cumulative then common stockholder received = 800000-240000 = 560000

When preferred stock is cumulative then common stockholder received = 800000-240000 = 560000